Starfax, Inc., manufactures a small part that is widely used in various electronic products such as home computers. Results for the first three years of operations were as follows (absorption costing basis):

Year 1 Year 2 Year 3
Sales $ 1,100,000 $ 838,000 $ 1,100,000
Cost of goods sold 860,000 608,000 910,000
Gross margin 240,000 230,000 190,000
Selling and administrative expenses 220,000 190,000 220,000
Net operating income (loss) $ 20,000 $ 40,000 $ (30,000 )

  

In the latter part of Year 2, a competitor went out of business and in the process dumped a large number of units on the market. As a result, Starfax’s sales dropped by 20% during Year 2 even though production increased during the year. Management had expected sales to remain constant at 50,000 units; the increased production was designed to provide the company with a buffer of protection against unexpected spurts in demand. By the start of Year 3, management could see that it had excess inventory and that spurts in demand were unlikely. To reduce the excessive inventories, Starfax cut back production during Year 3, as shown below:

Year 1 Year 2 Year 3
Production in units 50,000 60,000 40,000
Sales in units 50,000 40,000 50,000

Additional information about the company follows:

  1. The company’s plant is highly automated. Variable manufacturing expenses (direct materials, direct labor, and variable manufacturing overhead) total only $5.20 per unit, and fixed manufacturing overhead expenses total $600,000 per year.

  2. A new fixed manufacturing overhead rate is computed each year based that year's actual fixed manufacturing overhead costs divided by the actual number of units produced.

  3. Variable selling and administrative expenses were $3 per unit sold in each year. Fixed selling and administrative expenses totaled $70,000 per year.

  4. The company uses a FIFO inventory flow assumption. (FIFO means first-in first-out. In other words, it assumes that the oldest units in inventory are sold first.)

Starfax’s management can’t understand why profits doubled during Year 2 when sales dropped by 20% and why a loss was incurred during Year 3 when sales recovered to previous levels.

Required:

1. Prepare a contribution format variable costing income statement for each year.

2. Refer to the absorption costing income statements above.

a. Compute the unit product cost in each year under absorption costing. Show how much of this cost is variable and how much is fixed.

b. Reconcile the variable costing and absorption costing net operating income figures for each year.

5b. If Lean Production had been used during Year 2 and Year 3, what would the company’s net operating income (or loss) have been in each year under absorption costing?
Answer #1

1.

Year 1 Year 2 Year 3
Sales $      1,100,000 $       838,000 $1,100,000
Less : Variable Costs
Cost of Goods Sold $          260,000 =50000*5.2 $       208,000 =40000*5.2 $    260,000
Selling and administrative expenses $          150,000 =50000*3 $       120,000 =40000*3 $    150,000
Total Variable Expenses $          410,000 $       328,000 $    410,000
Contribution Margin $          690,000 $       510,000 $    690,000
Less : Fixed Costs
Manufacturing Overhead $          600,000 $       600,000 $    600,000
Selling and administrative expenses $            70,000 $         70,000 $      70,000
Total Fixed Costs $          670,000 $       670,000 $    670,000
Net Operating Income (Loss) $            20,000 $     -160,000 $      20,000

2.a.

Product Cost
Year 1 Year 2 Year 3
Variable manufacturing expenses $                 5.20 $              5.20 $           5.20
Fixed manufacturing overhead expenses $               12.00 =600000/50000 $            10.00 =600000/60000 $         15.00 =600000/40000
Total Product Cost per unit $               17.20 $            15.20 $         20.20

2.b.

Reconciliation
Year 1 Year 2 Year 3
Income as per Absorption Costing $            20,000 $     40,000 $     -30,000
Less : Fixed Overhead carried forward $   200,000 =20000*10 $    150,000 =10000*15
Add : Fixed Overhead brought forward $    200,000 =20000*10
Income as per Variable Costing $            20,000 $   160,000 $      20,000

3.
If lean production has been used, income would have same as under Variable costing since Production is matched with sales. Therefore Income for Year 2 and Year 3 would have been ($160000) & $20000 respectively